OpenStudy (anonymous):

Find an exact solution to the equation below so that 0 < theta < pi/2. 1=sqrt(3)tan(2theta) I know that I have to use the double angle formula for tan but I'm not positive what to do after that! Any help is appreciated!

OpenStudy (anonymous):

nah, cube the equation first

OpenStudy (anonymous):

So 1=3tan(2theta) 1/3=tan(2theta) Still running into the issue of double angle! (:

OpenStudy (anonymous):

1/3=(2tanx)/(1-tan^2x)

OpenStudy (anonymous):

Not sure what to do from there though.

OpenStudy (anonymous):

where did the sqrt go??

OpenStudy (anonymous):

Oh shoot haha I thought of it as a cube root.

OpenStudy (anonymous):

Why cube it?

OpenStudy (anonymous):

coz me toothought of it as a cube root.

OpenStudy (anonymous):

whats the question. can u use the equation tab n write it?

OpenStudy (anonymous):

\[1=\sqrt{3}\tan(2x)\]

OpenStudy (anonymous):

1=√3tan(2x) 1/√3 = tan(2x) 2x = 30 x = 15

OpenStudy (anonymous):

simple, lol

OpenStudy (anonymous):

How did you get from the second step to the third?

OpenStudy (anonymous):

and it should be in radians sorry. (theta)

OpenStudy (anonymous):

oh, then 2x = pi/6

OpenStudy (anonymous):

15 degrees = pi/12 radians which is correct thank you.

OpenStudy (anonymous):

welcome :D